Karlottalotta:

Are the following figures similar? (1 point) Kites ABCD and EFGH are shown. AB equals 2. BC equals 10. EF equals 5. FG equals 20. Yes; the corresponding angles are congruent. No; the corresponding angles are not congruent. Yes; the corresponding sides are proportional. No; the corresponding sides are not proportional.

Karlottalotta:

1 attachment
XioGonz:

Do you know how it looks like when corresponding angles are congruent or not?

XioGonz:

1 attachment
Karlottalotta:

yes

Karlottalotta:

I believe its C

XioGonz:

Why do you believe it's C?

XioGonz:

@jhonyy9 Do you agree with her?

jhonyy9:

you say that 2/10 = 5/20 ?

jhonyy9:

corresponding sides proportionality

jhonyy9:

so bc. 2/10 = 1/5 not equal 5/20=1/4 the last one choice is true

jhonyy9:

@XioGonz

Karlottalotta:

thank you